During open market operations the Federal Reserve Bank purchases $120 million dollars worth of securities. The estimate, using the simply multiplier model is that this will raise checkable deposits in the economy by $750 million. In this, the required reserve ratio is

Group of answer choices

A) 12%

B) 16%

C) 18%

D) 20%
Answer #1

Answer : The answer is option B.

Changes in money supply = Changes in deposits * Money multiplier

=> 120 = 750 * Money multiplier

=> Money multiplier = 120 / 750 = 0.16 * 100 = 16%

Therefore, option B is correct.